OpenStudy (anonymous):

For the limit, either evaluate immediately, or if it is indeterminate, find the form it represents, and then evaluate using the methods of this section. lim x → 0 (1 + tan(2x))^(3/x)

OpenStudy (zugzwang):

Well, it's indeterminate all right, just find which form?

OpenStudy (anonymous):

and then evaluate it that's the hard part

OpenStudy (anonymous):

I know it's 0^0

OpenStudy (zugzwang):

\[\lim_{x \rightarrow 0}\left( 1 + \tan2x \right)^{\frac{3}{x}}\] This is actually of the form \[1^{\infty}\]

OpenStudy (zugzwang):

1 + tan 2x evaluated at x = 0 is just 1 3/x goes to infinity as x goes to 0.

OpenStudy (zugzwang):

Can you do it from here?

OpenStudy (anonymous):

and when I get 1/0 I can't use LR so what do I do?

OpenStudy (zugzwang):

Suppose \[L = \lim_{x \rightarrow 0}\left( 1 + \tan2x \right)^{\frac{3}{x}}\]

OpenStudy (zugzwang):

Something doesn't want me to do this...

OpenStudy (anonymous):

What?

OpenStudy (zugzwang):

Hang on, I'll fix something up...

OpenStudy (zugzwang):

Ok, have a look at this